OpenStudy (anonymous):

What is the x-intercept of the line given by the equation 2x + 3y = 12

hartnn (hartnn):

to get x intercept, just put y=0 in your equation

hartnn (hartnn):

what u get if you put y= 0 in 2x+3y =12 ?

OpenStudy (anonymous):

2x - 3(0) = 12 2x - 3 = 12

hartnn (hartnn):

3*0 = 0

hartnn (hartnn):

2x = 12 x=...?

OpenStudy (anonymous):

2*6 = 12

hartnn (hartnn):

so, x= 6 yes :)

hartnn (hartnn):

x-intercept = (6,0)

OpenStudy (anonymous):

So y always equals 0?

hartnn (hartnn):

only for x-intercept, yes

OpenStudy (anonymous):

what about y-intercept? does that make x equal zero, or no?